OpenStudy (anonymous):

which function below is the inverse of f(x)=3x+5 7

OpenStudy (anonymous):

You have a function y = 3x + 5. To find the inverse, interchange x with y, and y with x, and then solve for y. Let's do that: We get x = 3y + 5 Now, let's solve for y, we get 3y = x - 5 so y = (x-5)/3 which is the inverse. The inverse of f(x), is symbolized by f^(-1)(x), so we can say f^(-1)(x) = (x-5)/3

OpenStudy (anonymous):

im sorry is 3x+5 over 7

OpenStudy (anonymous):

Do the problem identically to the way I just showed you above.

OpenStudy (anonymous):

(7x-5)/3